0.9m Telescope "Events"


____________________________________________________________________________

   DATE         EVENT                      RESULTS
____________________________________________________________________________

16 MAY 2016     Torrent controller replaces ARCON

09 FEB 2016     primary washed

17 OCT 2015     1.5m spec ARCON video board installed only ur amplifier works

19 AUG 2015     baffle re-installed and lateral support fixed
19 AUG 2015     primary washed

03 DEC 2014     baffle re-aligned - was out of position due to loose support 
                (from mirror cleaning?)

13 AUG 2014     primary washed

22 JAN 2014     first night of dewar & CCD replacement 

20 JAN 2014     dewar removed due to coolant leak & vacuum loss

                20 yr old O-ring replaced
                L2 tank temperature sensor wires replaced due to overheating
                molecular sieve & vacuum sensor changed
                detector cleaned

24 APR 2012     primary washed             86% reflectivity
24 APR 2012     measured                   63% reflectivity

30 DEC 2010     recabling                  rebalanced
		electronics repositioned

27 DEC 2010     primary washed

24 JUN 2010     primary washed             87% reflectivity
24 JUN 2010     measured                   75% reflectivity

11 SEP 2009     primary washed
                guider cleaned

07 MAR 2009     power outage               gain = 1 on CCD saturates ~40K

05 FEB 2009     primary washed             86% reflectivity
28 JAN 2009     measured                   48% reflectivity

25 APR 2005     first night of new TCS     pointing to < 20 arcsec, fast slew
02 APR 2005     last night of old TCS

   MAR 2005     Tek #1 V filter used       replaced Tek #2 for CTIOPI
   FEB 2005     Tek #2 V filter cracked   
____________________________________________________________________________

Mirror Washes

    https://www.ctio.noao.edu/DocDB/0002/000291/019/Mirror_Reflectance.pdf

Mirror Washing Schedule

    https://www.ctio.noao.edu/DocDB/0005/000597/001/Mirror%20Cleaning%20Schedule%202013.pdf